An electron with a speed of 6.79 × 108 cm/s in the positive direction of an x axis enters an electric field of magnitude 2.13 × 103 N/C, traveling along a field line in the direction that retards its motion. (a) How far will the electron travel in the field before stopping momentarily, and (b) how much time will have elapsed? (c) If the region containing the electric field is 5.57 mm long (too short for the electron to stop within it), what fraction of the electron’s initial kinetic energy will be lost in that region?
Answer #1

(a) acceleration =-eE/m=-3.75*10^(14) m/s^2

initial velocity =6.79*10^6 m/s

using

v^2=u^2+2aS

we get S=distance travelled before stopping=0.0615 m =6.15 cm

(b) time =u/a=1.8*10^(-8) sec

(c) velocity at S=5.57 mm

v^2=u^2+2aS

v^2=u^2+2as

v=6.47*10^(6) m/s

fraction of energy lost =(u^2-v^2)/u^2=0.09 fraction lost =9/100 fraction lost